OpenStudy (anonymous):

What is x^2+5x+4 factored?

OpenStudy (anonymous):

(X+1)(x+4)

OpenStudy (anonymous):

x2 - 5x + 4 = x2 - 4x - x + 4 = x(x - 4) - (x - 4) = (x - 4)(x - 1)

OpenStudy (zzr0ck3r):

\(1*4=4\) we want two numbers that add to \(5\) but multiply to \(4\) \(1,4\) are two such numbers expand the middle term with these numbers \(x^2+5x+4\rightarrow x^2+1x+4x+4=x^2+x+4x+4\) now factor what you can from the first two terms \(x(x+1)=4x+4\) now factor what you can from the second pair of terms to make it look \((x+1)\) \(x(x+1)+4(x+1)\) now factor out the \((x+1)\) \(x^2+5x+4=(x+1)(x+4)\)

OpenStudy (zzr0ck3r):

\(x(x+1)=4x+4\) should say \(x(x+1)+4x+4\)

OpenStudy (zzr0ck3r):

more concise \(\begin{equation} \begin{aligned} x^2+5x+4 &= x^2+x+4x+4 \\ &=x(x+1)+4(x+1) \\ &=(x+1)(x+4) \end{aligned} \end{equation}\)

OpenStudy (zzr0ck3r):

@Hapreet the rules say we are not here to just give out answers, as this will not help anyone..... please try and explain things. I know you see me typing and want to get it out fast, but this is not the point of the site.

OpenStudy (anonymous):

Oh well, I wasn't planning to response to this comment^, but it had been bothering me since yesterday. For all I know, I did see you and the piercethehorizon typing, but you could also see piercethehorizon typing. It seems like you were also competing to get the answer out there. More importantly, if you have seen the other questions that I have answered I always give the answer and then I explain how I got them. I don't always just give out the answers.

OpenStudy (anonymous):

One last thing I want to say is that thanks for explaining this answer before I did and I am sorry for blurping out the answer before you did